can you please help me understand how to simplify this problem (3+2i)(5-i)
i keep getting different answers each time i try it the first step that i did was to use the distributive property and then i got 15-3i+10i-2 then i simplified it and got 7i+13

Found 2 solutions by mszlmb, adesouza:
Answer by mszlmb(115)

SO CLOSE!!!

When you multiplied it out, you slyly worked out that i2=1 when it really equals -1!

So from (3+2i)(5-i) you foil to get 15-3i+10i-2i2

Since i2=-1, you can replace the i2 with -1 to obtain:
15-3i+10i-2(-1) which equals 15-3i+10i+2 which simplifies to 17+7i.

You've got the concept down though ☺

Answer by adesouza(1)

(3+2i)(5-i)
=15 - 3i + 10i - 2i^2
=15 + 7i - 2(-1) (Note: i^2 = -1)
=15 + 7i + 2
=7i + 17